### v1

# VesselSAM: Leveraging SAM for Aortic Vessel Segmentation with LoRA and Atrous Attention

Abstract
Medical image segmentation is crucial for clinical diagnosis and treatment planning, particularly for complex anatomical structures like vessels. In this work, we propose VesselSAM, a modified version of the Segmentation Anything Model (SAM), specifically designed for aortic vessel segmentation. VesselSAM incorporates AtrousLoRA, a novel module that combines Atrous Attention with Low-Rank Adaptation (LoRA), to improve segmentation performance. Atrous Attention enables the model to capture multi-scale contextual information, preserving both fine local details and broader global context. At the same time, LoRA facilitates efficient fine-tuning of the frozen SAM image encoder, reducing the number of trainable parameters and ensuring computational efficiency. We evaluate VesselSAM on two challenging datasets: the Aortic Vessel Tree (AVT) dataset and the Type-B Aortic Dissection (TBAD) dataset. VesselSAM achieves state-of-the-art performance with DSC scores of 93.50%, 93.25%, 93.02%, and 93.26% across multiple medical centers. Our results demonstrate that VesselSAM delivers high segmentation accuracy while significantly reducing computational overhead compared to existing large-scale models. This development paves the way for enhanced AI-based aortic vessel segmentation in clinical environments. The code and models will be released at https://github.com/Adnan-CAS/AtrousLora.
Index Terms:
SAM, parameter efficient fine-tuning, LoRA, atrous attention, aortic vessel segmentation
I Introduction
Medical imaging stands at the cutting edge of modern healthcare, serving a vital tool in diagnosing and treating various diseases. Within this domain, medical image segmentation is a critical component aiming to delineate structures such as organs, tumors, and vessels [1]. Aortic vessel segmentation is particularly important for diagnosing cardiovascular diseases, enabling precise assessments of vascular health and supporting interventions such as stent placement and aneurysm monitoring. It plays a crucial role in computer-aided diagnosis, treatment planning, and surgical interventions [2]. With the rapid advancement of computational resources and the growing availability of medical data, Vision Transformers (ViTs) have emerged as a transformative approach in medical image analysis [3]. Unlike traditional convolutional models, ViTs utilize self-attention mechanisms to capture long-range dependencies and global context [4], significantly improving the ability to model complex structures within medical images [5].
This paradigm shift has paved the way for more advanced segmentation techniques like the Segmentation Anything Model (SAM) [6], Swin-Unet [7], UNETR [8], SAMMedAI [9], and MedSAM [10] which leverage the power of ViT’s for accurate and efficient segmentation tasks. SAM enables users to generate segmentation masks using interactive prompts like clicks, bounding boxes and text. Its remarkable zero-shot and few-shot abilities have proven highly effective for natural images, gaining wide attention. However, while SAM excels in natural image segmentation, recent studies have highlighted its limitations in the medical domain[11][12].
Medical images, often characterized by low contrast, ambiguous tissue boundaries, and small regions of interest, present unique challenges to SAM [13]. Several recent approaches [14, 15, 16] have sought to fine-tune SAM for medical image segmentation by incorporating domain-specific enhancements. However, fine-tuning these models requires significant computational resources, due to the extensive number of parameters in foundation models like SAM. Furthermore, training large models with limited task-specific data often results in overfitting and inferior performance. To overcome these challenges, parameter-efficient fine-tuning (PEFT) approaches such as Low-Rank Adaptation (LoRA) [17] have been developed as a viable solution. Various methods have incorporated LoRA with SAM to address the computational efficiency while maintaining performance gains, specifically for medical image segmentation[18][19].
Despite these efforts, several fundamental intrinsic limitations of SAM persist. SAM’s image encoder, based on plain ViTs, inherently lacks crucial vision-specific inductive biases needed to capture local patterns and fine details essential for dense predictions in medical imaging [20]. Additionally, SAM’s ViT-based architecture relies on global attention without integrating regional attention or sparse attention mechanisms, which are vital for focusing on relevant regions and reducing computational overhead [21]. While regional attention, helps capture spatial hierarchies at different scales, SAM’s reliance on global attention limits its ability to focus on smaller, intricate regions in medical images. On the contrary, the absence of sparse attention prevents SAM from efficiently capturing global context without a significant computational cost. These limitations make SAM prone to errors, such as hallucinating small, disconnected components in its segmentation [4][9], especially when modeling structures like vessels, tumors, or lesions. To enhance the performance of plain ViTs in dense prediction tasks, recent research has combined Transformer and convolutional features [22] [23]. Work [24] integrates atrous attention with ViTs, enabling multi-scale feature extraction while preserving resolution. Atrous attention combines regional and sparse attention, allowing the model to focus on local details while still capturing the broader context.
Inspired by the work[24], we propose VesselSAM, a model that integrates Atrous Attention with SAM, leveraging both global transformer attention and local convolutional inductive biases. VesselSAM incorporates several key innovations to enhance SAM’s capabilities. First, we integrate Atrous Spatial Pyramid Pooling (ASPP) to capture multi-scale contextual information, enabling the model to handle both small and large anatomical structures without sacrificing spatial resolution[22]. Additionally, Atrous Attention mechanisms are introduced, combining dilated windows at different scales to balance local feature extraction with global contextual understanding, allowing the model to focus on fine details while maintaining a comprehensive view of the entire image [23]. Furthermore, VesselSAM incorporates LoRA[18] layers to fine-tune the model efficiently, reducing the need for computationally expensive full retraining while ensuring high performance across various medical segmentation tasks.
The main contributions of this work are summarized as follows:
  • We propose VesselSAM, a novel model that integrates AtrousLoRA,a module that combines Atrous Attention within Low-Rank Adaptation (LoRA) to enhance the SAM architecture for vascular image segmentation, particularly aortic vessel segmentation. AtrousLoRA enables VesselSAM to capture both local and global features efficiently, improving segmentation accuracy while keeping the pre-trained image encoder frozen. The integration of Atrous Attention allows for multi-scale feature extraction through dilated convolutions, while LoRA reduces the number of trainable parameters without compromising model performance.
  • AtrousLoRA is integrated into VesselSAM and comprises two key modules: the Atrous Spatial Pyramid Pooling (ASPP) module and the Attention mechanism. The ASPP module uses dilated convolutions at different rates to capture multi-scale contextual information, enabling VesselSAM to focus on both fine details, such as small vessel boundaries, and broader anatomical structures without losing spatial resolution. Meanwhile, the Attention mechanism balances local feature extraction with global context, guiding VesselSAM to focus on relevant anatomical regions and improving segmentation performance.
  • AtrousLoRA leverages LoRA’s key concept of applying a low-rank constraint to the transformer features, enabling VesselSAM to fine-tune the model efficiently with only 7% of the trainable parameters. This reduces the computational cost significantly, making VesselSAM more suitable for tasks with limited data, as it eliminates the need for full retraining while maintaining high performance.
  • We evaluate VesselSAM on multiple challenging benchmark datasets, including the Aortic Vessel Tree (AVT) Segmentation dataset and the imageTBAD dataset. The results show that VesselSAM, with AtrousLoRA, outperforms baseline methods in segmentation accuracy, robustness, and computational efficiency, particularly for aortic vessel segmentation.
The rest of this article is organized as follows. Section II provides an overview of related work, focusing on medical foundation models, parameter-efficient fine-tuning strategies, and the application of atrous convolution in Vision Transformers (ViTs). Section III details the methodology, beginning with a preliminary description of the SAM architecture, followed by the introduction of VesselSAM, LoRA, and AtrousLoRA, as well as the integration of the Atrous Attention Module for capturing both local and global features. Section III further discusses the role of the Prompt Encoder and Mask Decoder in generating high-resolution segmentation masks. Section IV outlines the experimental setup, including dataset descriptions, evaluation metrics, and presents both quantitative and qualitative results. It also includes an ablation study, addressing the model’s limitations and suggesting potential future directions. Finally, Section V concludes the article.
Figure 1: The architecture of the proposed VesselSAM framework. The model combines frozen pre-trained transformer blocks with learnable LoRA-Atrous Attention layers, enabling enhanced multi-scale feature extraction through Atrous Spatial Pyramid Pooling (ASPP) and attention mechanisms. Image embeddings generated by the frozen image encoder are fused with a bounding box prompt from the frozen prompt encoder and processed by the learnable mask decoder to produce the final segmentation. Blue represents frozen components, while orange denotes learnable parameters.
II Related Work
II-A ViT and SAM Based Medical Foundation Models
Vision Transformers (ViTs) based medical foundation models have significantly impacted medical image segmentation, with models like UNETR[8] leading the way. UNETR utilizes a ViT-based encoder to effectively capture global context while integrating it with a U-Net architecture for effective medical image segmentation. In contrast, SAM-based medical foundation models leveraging transformer architectures have demonstrated impressive performance across a broad range of natural image segmentation tasks. However, their application in the medical domain has faced limitations due to the unique challenges of medical images, such as low contrast and complex anatomical structures. Recognizing these limitations, MedSAM[10] sought to improve SAM’s performance in medical image segmentation by freezing the large pre-trained image encoder and prompt encoder, while fine-tuning only the lightweight mask decoder on domain-specific medical datasets. This approach leverages SAM’s powerful pre-trained architecture while adapting its mask prediction capabilities to the medical domain.
II-B Parameter-Efficient Model Fine-Tuning
The concept of Parameter-Efficient Fine-Tuning (PEFT) has emerged as an effective way to adapt large foundational models like SAM to specific downstream tasks with minimal additional parameter costs. One prominent PEFT approach, LoRA (Low-Rank Adaptation), has been successfully incorporated into SAM-based models. For example, SAMed[11] applied LoRA to SAM’s frozen image encoder, fine-tuning the LoRA layers, the prompt encoder, and the mask decoder together on medical datasets like Synapse multiorgan, demonstrating significant performance improvements. Similarly, SAMAdp[19] introduces a lightweight adapter module to enhance SAM’s performance in challenging segmentation tasks. By integrating task-specific prompts and adapters, it improves accuracy while maintaining computational efficiency, demonstrating adaptability across various domains. Other works have taken different approaches to enhancing SAM for medical applications. SAMMed[9] evaluated SAM across 53 public medical imaging datasets, revealing that while SAM has strong zero-shot segmentation capabilities but it often underperforms without fine-tuning.
II-C Atrous Convolution in ViTs
The use of Atrous Convolution (dilated convolution) in ViTs has gained attention as a powerful method to capture both local priors and global context, essential for segmentation tasks[21]. Atrous convolution increases the receptive field by ”skipping” pixels, allowing the model to capture information over larger regions without downsampling, thus preserving fine-grained details while enhancing the ability to model broader spatial relationships. This technique, initially popularized in DeepLab [6] for convolutional networks, has proven effective in extracting multi-scale features, which are crucial for complex segmentation tasks involving varying object sizes. In ViTs, where image features are typically processed as non-overlapping patches, integrating Atrous Convolutions enables the model to learn hierarchical spatial dependencies. By applying dilated convolutions at multiple rates, Atrous Spatial Pyramid Pooling (ASPP) modules allow the model to capture multi-scale contextual information, bridging the gap between local interactions and global dependencies. This approach is particularly beneficial in tasks requiring detailed segmentation, where capturing both local fine details and global context is necessary for accurate predictions. Recent advancements have shown that Atrous Convolutions are crucial for improving the performance of ViTs in segmentation tasks, particularly in domains such as medical imaging. In our model, we leverage the power of ASPP and Attention mechanisms to enhance the ViT encoder’s ability to capture both local priors and global context, effectively enabling the model to handle complex, high-resolution segmentation tasks with greater accuracy.
Figure 2: LoRA and AtrousLoRA: A comparative overview with detailed insights into the Atrous Attention Module. Both LoRA and Atrous-LoRA introduce a trainable encoder-decoder structure that operates in parallel with frozen pre-trained weights. (a) LoRA applies a low-rank constraint on the weight updates by factorizing them into smaller matrices (b) AtrousLoRA extends this approach by incorporating Atrous Attention Module into the bottleneck of LoRA, leveraging multi-scale dilated convolution operations for enhanced feature extraction. (c) The Atrous Attention Module features an Atrous Spatial Pyramid Pooling (ASPP) module with various dilation rates, global image-level pooling, and an attention mechanism that refines feature maps through element-wise multiplication with attention weights.
III Methodology
III-A Overview
VesselSAM is a promptable segmentation model designed to enhance vascular structure segmentation in medical images. Building upon the standard SAM (Segment Anything Model) framework, VesselSAM integrates key modifications such as the Atrous Attention Module and LoRA (Low-Rank Adaptation) layers to improve performance for medical image segmentation. The image encoder and prompt encoder from the SAM model are frozen to preserve their pre-trained features, while the Atrous Attention module and LoRA layers enhance the model’s ability to capture multi-scale features and optimize training efficiency. The final segmentation is generated by a mask decoder that refines the fused embeddings using cross-attention mechanisms. The overall design ensures VesselSAM is a robust and efficient model for medical image segmentation tasks, particularly vascular segmentation in aortic imaging.
III-B Preliminary: SAM architecture
The SAM[6] is a prompt-based segmentation framework composed of three main components: the Image Encoder, Prompt Encoder, and Mask Decoder. The Image Encoder is based on a ViT, which processes input images using 16 × 16 pixel patches through transformer blocks to capture image features, resulting in image embedding. The Prompt Encoder handles various prompts, including points, bounding boxes and masks, converting them into feature vectors that guide the segmentation. The Mask Decoder is a two-layer transformer-based decoder that fuses image embedding and prompt features using cross-attention. It incorporates a Multi-Layer Perceptron (MLP) for feature refinement and dimensionality alignment, and utilizes convolutional layers for upsampling to produce high-resolution masks.
III-C VesselSAM
The VesselSAM architecture builds on the foundation of SAM, with several key modifications aimed at improving vascular structure segmentation in medical images. As illustrated in Fig. 1, VesselSAM incorporates the Atrous Attention module and LoRA layers, designed to capture multi-scale features and reduce the number of trainable parameters while maintaining segmentation accuracy.
In this design, the image encoder and prompt encoder from the original SAM architecture are frozen to retain their powerful pre-trained features. The image encoder, based on a Vision Transformer (ViT), extracts rich visual features from the input medical images. The prompt encoder processes sparse prompts such as points or bounding boxes, which guide the segmentation process by focusing on specific regions of interest in the image.
To enhance the model’s ability to capture both local and global features, the Atrous Attention module is integrated into the frozen image encoder. This module utilizes dilated convolutions to expand the receptive field, allowing the model to capture multi-scale features, which are crucial for medical images like small tumors or vascular boundaries.
Additionally, LoRA (Low-Rank Adaptation) layers are inserted between the transformer blocks in the image encoder. These layers compress the transformer features into a low-rank space and then re-project them, allowing efficient adaptation of the features while preserving the frozen transformer parameters. This modification improves training efficiency, reducing the number of trainable parameters and enhancing the model’s performance with fewer resources.
The final segmentation is generated by the mask decoder, which consists of a lightweight transformer decoder and a segmentation head. During training, the mask decoder is fine-tuned to refine the fused embeddings from the image and prompt encoders using cross-attention mechanisms. This ensures that the model is able to accurately segment fine-grained details, such as vascular structures, while also preserving broader anatomical context.
III-D LoRA and AtrousLoRA
LoRA [17] has emerged as a PEFT method, enabling task-specific adaptations of pre-trained models while significantly reducing computational and memory overhead. LoRA introduces low-rank trainable matrices to approximate weight updates, effectively bypassing the need to fine-tune the entire model(Fig. 2(a)). Instead, it adds two small matrices, $W_{e}$ and $W_{d}$, while keeping the original weights $W_{0}$ frozen during training. Given a pre-trained weight matrix $W_{0}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}}$, LoRA modifies the forward pass of the model as
$h=W_{0}x+W_{d}W_{e}x,$ (1)
where $W_{0}$ is the frozen pre-trained weight matrix, $W_{e}\in\mathbb{R}^{r\times C_{\text{in}}}$ and $W_{d}\in\mathbb{R}^{C_{\text{out}}\times r}$ are the low-rank encoder and decoder matrices, and $r$ is the rank of the decomposition, with $r\ll\min(C_{\text{in}},C_{\text{out}})$. Here, $x\in\mathbb{R}^{B\times C_{\text{in}}}$ represents the input, where $B$ is the batch size.
While LoRA is highly efficient for adapting pre-trained models, it lacks the ability to explicitly capture multi-scale contextual information, which is critical for vision tasks such as image segmentation and dense prediction. To address this limitation, we introduced Atrous LoRA which incorporates atrous (dilated) convolutions into the LoRA framework(Fig. 2(b)). Atrous convolutions expand the receptive field of the model without increasing the number of parameters, enabling it to capture both local and global dependencies.
Mathematically, with AtrousLoRA, Equation 1 changes to:
$h=W_{0}x+W_{d}\cdot\text{Atrous}(W_{e}x),$ (2)
where $W_{0}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}}$ is the frozen pre-trained weight matrix, $W_{e}\in\mathbb{R}^{r\times C_{\text{in}}}$ and $W_{d}\in\mathbb{R}^{C_{\text{out}}\times r}$ are the low-rank encoder and decoder matrices, and $x\in\mathbb{R}^{B\times C_{\text{in}}\times H\times W}$ is the input feature map. In this case, $B$ represents the batch size, $C_{\text{in}}$ and $C_{\text{out}}$ are the input and output channels, and $H$ and $W$ represent the height and width of the feature maps. The Atrous module applies atrous convolutions with predefined dilation rates to $W_{e}x$, effectively capturing multi-scale contextual features. Atrous LoRA employs fixed dilation rates, which simplifies implementation while maintaining adaptability for various vision-related tasks.
AtrousLoRA retains the efficiency of LoRA while extending its applicability to tasks requiring spatial and contextual understanding, such as semantic segmentation and medical image analysis. The predefined dilation rates ensure a balance between computational efficiency and multi-scale feature extraction.
Algorithm 1 Atrous Attention Module
  1: Input: $X\in\mathbb{R}^{B\times C\times H\times W}$
  2: Output: $Y_{\text{out}}\in\mathbb{R}^{B\times C\times H^{\prime}\times W^{\prime}}$
  3: $Y_{i}=f_{\text{dil}}(X;W_{i},d_{i}),\quad i=1\dots n$
  4: $Z=\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}X_{b,c,h,w}$
  5: $Y=[Y_{1},Y_{2},\dots,Y_{n},Z]$
  6: $Y_{\text{ASPP}}=\text{ReLU}(\text{BN}(f_{\text{1x1}}(Y)))$
  7: $A=\sigma(f_{\text{1x1}}(Y_{\text{ASPP}}))$
  8: $Y_{\text{out}}=Y_{\text{ASPP}}\odot A$
III-E Atrous Attention Module
We propose a novel attention mechanism for vision transformers, called Atrous Attention Module (Fig. 2(c)), which performs a fusion between regional and sparse attention. This approach allows us to capture both global context and local detail with efficient computational complexity, while preserving the hierarchical information present in medical images. Inspired by atrous convolution [24], which increases the receptive field by skipping rows and columns in the input feature map without adding extra parameters, Atrous Attention enables VesselSAM to focus on relevant anatomical structures across multiple scales. The process is shown in Algorithm 1.
The data flow within the Atrous Attention Module starts by passing the input feature map $X\in\mathbb{R}^{B\times C\times H\times W}$ through the Atrous Spatial Pyramid Pooling (ASPP) [21], which applies dilated convolutions at different rates $d_{i}$ to capture features at various scales. Each atrous convolution produces an output feature map $Y_{i}=f(X;W_{i},d_{i})$ where $W_{i}$ are the convolution weights and $d_{i}$ is the dilation rate. Additionally, global average pooling is performed on $X$ to obtain $Z=f_{\text{1x1}}(\text{GAP}(X))$ The outputs from ASPP, including the atrous convolutions $Y_{i}$ and the global pooling result $Z$, are concatenated into a Concatenated Feature Map $Y=[Y_{1},Y_{2},\dots,Y_{n},Z]$. This concatenated feature map then goes through a 1x1 Convolution, reducing it to the desired number of output channels, followed by Batch Normalization (BN) and ReLU activation, generating the ASPP Output $Y_{\text{ASPP}}=\text{ReLU}(\text{BN}(f_{\text{1x1}}(Y)))$.
This output is further processed through another 1x1 Convolution to create the Attention Map $A=f_{\text{1x1}}(Y_{\text{ASPP}})$ where $A\in\mathbb{R}^{B\times 1\times H\times W}$ is the attention map and a Sigmoid activation is applied to obtain $A_{\text{sigmoid}}=\sigma(A)$ which constrains the attention values between 0 and 1 that is where $A_{\text{sigmoid}}\in[0,1]^{B\times 1\times H\times W}$. Finally, the ASPP Output is multiplied element-wise with the Attention Map, producing the Final Output $Y_{\text{out}}=Y_{\text{ASPP}}\odot A_{\text{sigmoid}},$ The result is an attention-weighted feature map $Y_{\text{out}}\in\mathbb{R}^{B\times C^{\prime}\times H\times W}$, where important regions of the feature map are enhanced, and less important regions are suppressed. This mechanism enhances VesselSAM’s ability to focus on the most important features, improving segmentation accuracy while maintaining context from multiple scales.
III-F Atrous Spatial Pyramid Pooling
Atrous Spatial Pyramid Pooling (ASPP) is another important component in the VesselSAM Model. It plays a crucial role in enhancing the model’s ability to segment vascular structures by capturing multi-scale contextual information from medical images. VesselSAM, which integrates advanced segmentation techniques, utilizes ASPP to improve the segmentation of blood vessels by applying dilated convolutions with varying dilation rates. This allows the model to capture both fine details and broader contextual information without losing resolution. ASPP increases the receptive field by using dilated convolutions at multiple rates, enabling the model to understand both local features of vessels and their spatial relationships within the image, which is critical for accurate vascular segmentation.
Mathematically, in VesselSAM, ASPP begins by applying dilated convolutions with different dilation rates $d_{i}\in\{d_{1},d_{2},\dots,d_{n}\}$, where each rate captures features at different scales. For each dilation rate $d_{i}$, the dilated convolution operation is applied to the input feature map $X$ as follows:
$Y_{i}=f_{\text{dil}}(X;W_{i},d_{i})=X*_{d_{i}}W_{i},$ (3)
where $*_{d_{i}}$ denotes the dilated convolution, and $W_{i}$ represents the filter with dilation rate $d_{i}$. In addition to the dilated convolutions, a global average pooling operation is applied to capture the global context of the input image, which is mathematically defined as:
$Z=\frac{1}{H\times W}\sum_{h=1}^{H}\sum_{w=1}^{W}X_{b,c,h,w},$ (4)
where $Z$ represents the global pooling result, effectively reducing the spatial dimensions to $1\times 1$ while maintaining the channel information. These outputs are then concatenated into a single feature map:
$Y_{\text{concat}}=[Y_{1},Y_{2},\dots,Y_{n},Z],$ (5)
The concatenated feature map $Y_{\text{concat}}$ contains multi-scale information, helping VesselSAM capture both local vessel features and broader contextual relationships, which is essential for accurately segmenting complex vascular structures. To reduce the dimensionality of this concatenated feature map, a $1\times 1$ convolution is applied:
$Y_{\text{ASPP}}=f_{\text{1x1}}(Y_{\text{concat}})=W_{\text{1x1}}\cdot Y_{\text{concat}}+b_{\text{1x1}},$ (6)
where $W_{\text{1x1}}$ and $b_{\text{1x1}}$ are the weight and bias of the $1\times 1$ convolution. Finally, a non-linear activation function, such as ReLU, is applied to introduce non-linearity into the model:
$Y_{\text{ASPP}}=\text{ReLU}(Y_{\text{ASPP}})$ (7)
In VesselSAM, ASPP is integral to capturing multi-scale contextual features, allowing the model to handle objects, such as blood vessels, at different scales effectively. By using dilated convolutions with various rates, ASPP enables VesselSAM to segment vascular structures accurately while maintaining computational efficiency. This multi-scale feature extraction is essential for addressing the variability in vessel sizes and complex vascular networks commonly encountered in medical imaging.
III-G Prompt Encoder And Mask Decoder
In VesselSAM, the Prompt Encoder remains frozen, ensuring the stability of the pre-trained parameters while allowing for efficient processing of user prompts. In our case, the prompts are provided in the form of bounding boxes, which are represented by their top-left and bottom-right corner points. Each corner point is mapped into a 256-dimensional embedding, which serves as the input to the segmentation process. By freezing the prompt encoder, VesselSAM enables real-time interaction, as the image embedding can be precomputed, allowing the user to provide bounding-box input dynamically without the need for retraining.
On the other hand, the Mask Decoder in VesselSAM is fully trainable and plays a crucial role in producing the segmentation output. The decoder architecture includes two transformer layers, which are responsible for fusing the image embedding with the prompt embeddings through cross-attention. This fusion allows the bounding box information to guide the segmentation task effectively. Following this, the decoder employs two transposed convolution layers to upsample the combined embedding to a resolution of 256 × 256, ensuring a high level of detail is retained in the final segmentation mask. The output is then passed through a sigmoid activation function, followed by bi-linear interpolation to match the resolution of the original input image, thereby producing the final high-resolution mask.
TABLE I: Performance Comparison of our Proposed VesselSAM with other ViTBased and SAMbased Models on AVT dataset
Big Model
Methods #Params (M) /Ratio (%) AVT-Dataset Dongyang Hospital AVT-Dataset Rider Hospital AVT-Dataset Kits Hospital
DSC(%) $\uparrow$ HD(mm) $\downarrow$ DSC(%) $\uparrow$ HD(mm) $\downarrow$ DSC(%) $\uparrow$ HD(mm) $\downarrow$
UNET [25] 29.9 / 100 88.95 4.24 87.70 4.40 88.03 4.38
UNETR [8] 92.5 / 100 89.38 4.15 88.04 4.39 88.69 4.28
SAM-ViTb [6] 91.0 / 100 81.12 9.85 79.93 10.20 80.50 10.00
MedGIFT [26] 120.7 / 100 88.70 4.27 87.50 4.41 87.09 4.37
MedSAM-Vanilla [10] 93.7 / 100 89.50 4.13 87.04 4.46 88.65 4.27
MedSAM-FT [10] 93.7 / 100 92.49 3.64 90.35 4.01 91.45 3.95
SAMMed-Vanilla [9] 91.0 / 100 88.02 4.37 87.30 4.43 87.20 4.45
SAMMed-FT [9] 91.0 / 100 89.76 4.10 88.25 4.32 88.75 4.28
Small Model
SAMed-FT [11] 6.3 / 6.7 88.23 4.34 89.45 4.14 88.80 4.24
SAMAdp-FT [19] 4.1 / 4.3 90.30 4.02 89.75 4.10 89.90 4.05
VesselSAM 6.8 / 7.2 93.50 3.56 93.25 3.59 93.02 3.64
Note: Bold indicates the best results and underline denotes the second best results.
TABLE II: Performance Comparison of our Proposed VesselSAM with other ViTBased and SAMbased Models on TBAD Dataset
Big Model
Methods #Parms(M) / Ratio(%) DSC(%) $\uparrow$ HD(mm) $\downarrow$
UNET [25] 29.9 / 100 88.65 4.29
UNETR [8] 92.5 / 100 89.20 4.18
SAM-VitB [6] 91.0 / 100 79.53 10.15
MedGIFT [26] 120.7 / 100 87.60 4.49
MedSAM-Vanilla [10] 93.7 / 100 88.40 4.29
MedSAM-FT [10] 93.7 / 100 92.20 3.63
SAMMed-Vanilla [9] 91.0 / 100 89.40 4.14
SAMMed-FT [9] 91.0 / 100 87.40 4.40
Small Model
SAMed-FT [11] 6.3 / 6.7 88.20 4.43
SAMAdp-FT [19] 4.1 / 4.3 89.71 4.14
VesselSAM 6.8 / 7.2 93.26 3.58
Note: Bold indicates the best results and underline denotes the second best results.
IV Experiments
IV-A Datasets
In our experiments, we utilized two key datasets to evaluate the effectiveness of the proposed VesselSAM model in complex medical segmentation tasks. The Aortic Vessel Tree (AVT) Segmentation dataset [27] comprises 56 contrast-enhanced CT angiography (CTA) scans collected from three sources: the KiTS Grand Challenge, the Rider Lung CT dataset, and Dongyang Hospital. Among these, 38 cases were designated for training, while the remaining 18 were used for testing. All slices were resampled to a spatial resolution of 1 mm × 1 mm, with Hounsfield Unit (HU) values normalized to [0, 1]. Additionally, the TBAD dataset [28], comprising 100 CTA images from Guangdong Provincial People’s Hospital, was utilized for segmenting True Lumen (TL), False Lumen (FL), and False Lumen Thrombus (FLT) in Type-B Aortic Dissection (TBAD) cases. To conform to Segment Anything Model (SAM) requirements, both the AVT and TBAD datasets were converted from 3D CTA volumes into 2D slices. Each 3D scan was converted into NumPy arrays, with all slices resampled to a uniform resolution of 1 mm × 1 mm. Voxel intensity values were normalized using standard CT window settings [400, 40]. Ground truth masks were refined by removing labels of irrelevant structures and small objects, using thresholds of 1000 voxels for 3D volumes and 100 pixels for individual 2D slices. Only non-zero slices were retained, and intensity normalization was applied. The processed 2D slices were then resized to 1024 × 1024 pixels and converted into three-channel images by duplicating the grayscale slice across three channels (1024 × 1024 × 3), ensuring compatibility with SAM’s input format.
IV-B Loss Function and Evaluation Metrics
We have used a combined loss function comprising an unweighted sum between cross-entropy loss and Dice loss, which has been widely adopted for its robustness in medical image segmentation tasks[10]. Let $P$ represent the predicted segmentation output, and $T$ denote the corresponding ground truth. For each voxel $j$, $p_{j}$ and $t_{j}$ correspond to the predicted and ground truth values, respectively. The total number of voxels in the image is denoted by $M$. The binary cross-entropy loss is defined as:
$L_{CE}=-\frac{1}{M}\sum_{j=1}^{M}[t_{j}\log p_{j}+(1-t_{j})\log(1-p_{j})],$ (8)
where $L_{CE}$ quantifies the pixel-wise classification accuracy. The Dice loss, which measures the overlap between the predicted and ground truth regions, is given by:
$L_{D}=1-\frac{2\sum_{j=1}^{M}t_{j}p_{j}}{\sum_{j=1}^{M}(t_{j})^{2}+\sum_{j=1}^{M}(p_{j})^{2}},$ (9)
The final loss $L$ is computed as the sum of the cross-entropy loss and the Dice loss:
$L=L_{CE}+L_{D}$ (10)
This combined loss function ensures effective training by balancing region-based overlap and pixel-wise classification accuracy, making it suitable for a wide range of medical image segmentation tasks.
To evaluate the performance of the segmentation model, we employed two widely used metrics: Dice Similarity Coefficient (DSC) and Hausdorff Distance (HD). The DSC measures the spatial overlap between the predicted segmentation $P$ and the ground truth $T$, and is defined as:
$DSC(P,T)=\frac{2|P\cap T|}{|P|+|T|}$ (11)
where $|P\cap T|$ represents the intersection of the predicted and ground truth regions, and $|P|$ and $|T|$ denote the sizes of the predicted and ground truth regions, respectively. A higher DSC value indicates better segmentation accuracy, with a maximum value of 1 indicating perfect overlap.
The Hausdorff Distance (HD) quantifies the maximum distance between the boundaries of the predicted segmentation and the ground truth. It is defined as:
$HD(P,T)=\max\left(\sup_{x\in\partial P}\inf_{y\in\partial T}d(x,y),\sup_{y\in\partial T}\inf_{x\in\partial P}d(x,y)\right)$ (12)
where $\partial P$ and $\partial T$ represent the boundaries of the predicted and ground truth regions, respectively, and $d(x,y)$ is the Euclidean distance between points $x$ and $y$. A lower HD value indicates better boundary alignment between the predicted and ground truth segmentations.
These evaluation metrics provide complementary insights into the model’s performance, with DSC focusing on region-based overlap and HD emphasizing boundary accuracy. Together, they offer a comprehensive assessment of the segmentation quality.
Figure 3: Qualitative visual results on the AVT (Dongyang, Rider, Kits) and TBAD Datasets under Bounding Box Prompts. The first column shows the input images, followed by the ground truth (GT) in the second column. The subsequent columns show the results of various segmentation models: VesselSAM, SAM, MedSAM, SAMAdp, SAMedAdp, and SAM-MedIA. Each model’s output is overlaid with color-coded regions for true lumen (pink), false lumen (green), and the GT boundary line (yellow). The blue box represents the bounding box-prompt used for segmentation. The images have been zoomed in to enhance visibility.
IV-C Quantitative results
In this section, we provide a comprehensive comparison of our proposed method, VesselSAM, against various state-of-the-art (SOTA) models, including UNET[25], UNETR[8], SAM[6], MedSAM[10], SAMMed[9], SAMed[11] and SAM-Adopter[19]. Each method was assessed under the same conditions to ensure a fair comparison, allowing us to accurately evaluate performance metrics such DSC and HD. The results highlight how our approach outperforms the SOTA models in the field, demonstrating its effectiveness in tackling complex medical image segmentation tasks.
IV-C1 Quantitative Evaluation Results for AVT Dataset
The performance metrics for various segmentation methods on the Aortic Vessel Tree (AVT) dataset are presented in Table. I. This comparison encompasses both big and small models, illustrating the effectiveness of each approach across multiple hospitals. VesselSAM demonstrates exceptional segmentation performance, achieving a Dice Similarity Coefficient (DSC) of 93.50% at Dongyang Hospital, 93.25% at Rider Hospital, and 93.02% at Kits Hospital. This performance significantly surpasses that of state-of-the-art methods, including MedSAM and SAMAdp.
The incorporation of Atrous Attention and LoRA mechanisms within VesselSAM has contributed to its high performance, enabling the model to effectively capture multi-scale features essential for precise segmentation in medical imaging. In contrast, the other SAM-based models such as SAMAdp and SAMedAdp struggle with accuracy, as they produce a significant number of false positive regions within their segmentations. This disparity underscores the advantages of VesselSAM in accurately delineating vascular structures amidst challenging imaging contexts, ultimately supporting its utility for clinical applications.
IV-C2 Quantitative Evaluation Results for TBAD Dataset
The results for the Type-B Aortic Dissection (TBAD) dataset are summarized in Table. II, further highlighting the performance of VesselSAM. The model achieves a DSC of 93.26%, outperforming various competing methods, including UNETR and MedSAM. These findings illustrate VesselSAM’s robustness in accurately segmenting the true lumen (TL) and false lumen (FL), emphasizing its effectiveness in handling complex segmentation tasks within clinical settings.
In comparison, SAM and MedSAM display lower performance, with DSC scores of 79.53% and 92.20%, respectively. Moreover, other models such as SAMAdp and SAMedAdp also exhibit challenges in segmentation accuracy, as evidenced by their lower DSC values. The consistent high performance of VesselSAM across both the AVT and TBAD datasets demonstrates its potential as a valuable tool for medical image segmentation, particularly in complex cases where precision is paramount.
IV-D Qualitative results
To provide a more intuitive comparison, we present the qualitative segmentation results for various models, including VesselSAM, SAM, MedSAM, SAMAdp, SAMedAdp, and SAM-MedIA, as illustrated in the Fig. 3. The top row displays the results for aortic vessel segmentation, while the bottom row highlights the segmentation of true lumen (TL) and false lumen (FL) for Type-B Aortic Dissection (TBAD). In the aortic vessel segmentation task, VesselSAM effectively delineates the vessel structures, capturing intricate details that may be overlooked by other models. The segmentation accurately follows the boundaries of the aorta, demonstrating its robustness in identifying the vessel amidst surrounding tissues. In contrast, SAM struggles with segmentation accuracy, leading to significant misalignments with the ground truth, particularly in the definition of vessel edges. MedSAM demonstrates improved performance compared to SAM but still fails to capture some fine details, resulting in inaccuracies in the vessel’s contour. The models SAMAdp, SAMedAdp, and SAM-MedIA, struggles to accurately capture the true positive vessel areas, resulting in a significant number of false positive regions in their segmentations. While these models provide reasonable outputs, they tend to misidentify surrounding areas as part of the vessel structure.
In the segmentation of TL, FL and FLT in TBAD dataset, VesselSAM continues to excel by accurately capturing the luminal structures. The segmentation closely aligns with the GT, effectively distinguishing the TL, FL and the FLT. For better visualization, only the TL and FL are presented. In contrast, SAM encounters significant challenges, with poor segmentation of the TL, resulting in structural misrepresentations. MedSAM provides an improvement over SAM, but it still exhibits inaccuracies that affect its reliability in clinical applications. Other methods like SAMAdp, SAMedAdp, and SAM-MedIA similarly face challenges in accurately delineating the lumens, with occasional missing segments and imprecise boundaries.
Figure 4: Qualitative visual results on the AVT (Dongyang, Rider, Kits) and TBAD datasets using Bounding Box Prompts. The first column shows the input images, followed by the ground truth (GT) in the second column. Subsequent columns shows the results of the following configurations: ”VesselSAM w/o AMM” refers to the VesselSAM model with the MedSAM backbone but without the Atrous Attention Module (AMM), while ”VesselSAM* w/o AMM” uses the SAM backbone without AMM. ”VesselSAM* with AMM” and ”VesselSAM** with AMM” incorporate AMM with the MedSAM and SAM backbones, respectively.The outputs are color-coded to highlight the true lumen (pink), the GT boundary line (yellow) and the bounding box-prompt (blue).
TABLE III: Ablation study of Atrous Attention Module
Dataset VesselSAM* VesselSAM** Atrous Attention Module Dice Score
88.43%
89.56%
AVT-Dongyang [27] 91.23%
93.50%
88.25%
89.16%
AVT-KiTs[27] 91.57%
93.02%
87.89%
88.75%
AVT-Rider[27] 91.42%
93.25%
90.76%
91.68%
TBAD[28] 91.90%
93.26%
Note: VesselSAM* represents the VesselSAM Model with the MedSAM Model as a backbone and VesselSAM** represents the VesselSAM Model with the SAM Model as a backbone.
0010101010202020203030303040404040505050500.0020.0040.0060.0080.0100.012Converge faster andbetter!EpochsTraining Lossw Atrous Attention w/o Atrous Attention
Figure 5: The impact of Atrous Attention Module on validation loss over training epochs.
Figure 6: DSC VS Rank: Comparison of Dice Similarity Coefficients (DSC) for Aortic Vascular Tree (AVT) datasets (Dongyang, KiTs, Rider) and TBAD dataset across different LoRA ranks (2, 4, 16, 32, and 64), illustrating the performance stability and optimal rank selection for segmentation tasks.
IV-E Ablation Study
In this ablation study, we aimed to evaluate the effectiveness of different configurations of the VesselSAM (Segment Anything Model) on medical image segmentation tasks, particularly focusing on vessel segmentation. We conducted a series of comprehensive ablation experiments to evaluate the impact of key components in the model, including the Atrous Attention Module and the LoRA rank. These experiments help in understanding how the individual elements contribute to the overall performance of the model in terms of segmentation accuracy, using the Dice score as the primary evaluation metric.
In this ablation study, we aimed to evaluate the effectiveness of different configurations of the Vessel SAM (Segment Anything Model) on medical image segmentation tasks, particularly focusing on vessel segmentation. We conducted a series of comprehensive ablation experiments to evaluate the impact of key components in the model, Firstly compared the performance of two baseline models—Vessel SAM initialized with the MedSAM (medical domain-specific) and SAM (general domain) configurations. Secondly also tested an enhanced model incorporating Low-Rank Adaptation (LoRA) and Atrous Convolution. The objective was to analyze the impact of these variations on segmentation performance, using the Dice score as the primary evaluation metric.
IV-E1 Impact of the Backbone and Atrous Attention Module
In this ablation study, we assess the impact of the backbone architecture and the integration of the Atrous Attention module on the performance of VesselSAM. We compare two configurations: VesselSAM initialized with the MedSAM backbone (VesselSAM*) and the SAM backbone (VesselSAM**), which serve as the baseline models for this analysis. Additionally, we introduce the Atrous Attention module into both configurations to evaluate its effect on segmentation performance.
The Atrous Attention module is integrated into the image encoder to improve the model’s ability to capture multi-scale features. By utilizing dilated convolutions, this module expands the receptive field, enabling the model to focus on both small and large structures within the input image. This is particularly important for accurately segmenting vascular structures, where both fine details and broader contextual information are essential.
From the results presented in Fig. 4, it is evident that the Atrous Attention module improves the segmentation accuracy of both the MedSAM and SAM backbones. The segmentation outputs, which highlight the true lumen (pink), the GT boundary line (yellow), and the bounding box prompt (blue), demonstrate enhanced delineation of vascular structures when the Atrous Attention module is applied.
The quantitative results in Table III provide strong evidence supporting the effectiveness of integrating the Atrous Attention module with the MedSAM backbone. The configuration combining the MedSAM backbone with the Atrous Attention module (VesselSAM* with AAM) achieved the highest Dice score of 93.50% on the AVT-Dongyang dataset, outperforming all other configurations. This result highlights the significant benefit of using the MedSAM backbone, specifically designed for medical imaging, in combination with the Atrous Attention module, which enhances the model’s ability to capture multi-scale features. This combination provides a substantial improvement in segmentation accuracy, making it the most effective configuration for vascular segmentation.
In comparison, the VesselSAM model with the SAM backbone (VesselSAM** with AAM) also benefits from the Atrous Attention module, but the Dice scores are consistently lower. While these results still reflect an improvement over the baseline model with the Atrous Attention module, they demonstrate that the MedSAM backbone—tailored for medical applications—offers a clear advantage when combined with the Atrous Attention module.
These findings suggest that the Atrous Attention module consistently improves segmentation performance, but its full potential is realized when paired with a domain-specific backbone like MedSAM. This combination enables VesselSAM to achieve the best performance across multiple datasets, reinforcing the importance of both the backbone architecture and attention mechanism in improving segmentation accuracy.
The training dynamics are further illustrated in Fig. 6, where the training loss curves for both configurations are compared. The model with Atrous Attention (red line) shows faster convergence and lower validation loss compared to the model without the Atrous Attention module (green line). By around epoch 20, the model with Atrous Attention stabilizes at a lower training loss, indicating that the module accelerates convergence and enhances the model’s ability to segment vascular structures more accurately.
IV-E2 Impact of LoRA Rank
In this experiment, we investigated the effect of LoRA rank on the performance of VesselSAM. Low-Rank Adaptation (LoRA) is designed to reduce the number of trainable parameters, making the training process more efficient without compromising model performance. We tested different LoRA ranks (2, 4, 16, 32, and 64) and measured their impact on segmentation accuracy using the Dice score as the evaluation metric.
As illustrated in Fig. 6, the performance of VesselSAM showed significant variation across different LoRA ranks. LoRA rank 4 yielded the best performance, with the model achieving a Dice score of 93.5 on the AVT-Dongyang dataset, and similar strong performance on other datasets: 93.02 on AVT-KiTs, 93.25 on AVT-Rider, and 93.26 on TBAD. This suggests that LoRA rank 4 offers the optimal trade-off between segmentation accuracy and computational efficiency.
However, as the LoRA rank increased beyond 4, performance started to decline. For instance, at LoRA rank 16, the AVT-Dongyang Dice score dropped to 82.88, and at LoRA rank 32, it further decreased to 85.57. Interestingly, LoRA rank 64 resulted in slightly improved scores compared to rank 32, but still did not outperform rank 4. This trend indicates diminishing returns as the LoRA rank increases beyond an optimal point, with rank 4 providing the best overall segmentation performance.
IV-F Limitations and Future Work
In summary, this study shows that domain-specific models (MedSAM) combined with adaptation techniques such as LoRA and Atrous Convolution can outperform general-domain baselines, achieving superior segmentation accuracy. These insights are valuable for optimizing medical image segmentation models, particularly when faced with computational limitations. While VesselSAM demonstrates strong performance in vascular segmentation, there are a few limitations that need to be addressed. One key limitation is the reliance on bounding box prompts, which may not always provide enough detail for complex or ambiguous structures. To improve flexibility and accuracy, future work will incorporate alternative prompt mechanisms, such as text-based prompts, to provide richer, more intuitive guidance for segmentation tasks. Another limitation is the model’s dependence on high-quality input images. Although VesselSAM works well on clean, well-annotated datasets, its performance may decrease with noisy or low-resolution images. Future research will focus on enhancing the model’s robustness by incorporating data augmentation techniques and improving its generalization to diverse real-world imaging conditions.In addition, the integration of visual language models with VesselSAM offers exciting potential for future work. By combining the power of language understanding with visual information, these models could further enhance prompt generation and segmentation accuracy, enabling the model to better handle ambiguous or novel vascular structures with minimal user input. Furthermore, while VesselSAM is currently focused on segmentation of the aortic vessels, its potential for other vascular structures and medical domains has not been fully explored. Expanding the model’s applicability to other areas, such as brain or coronary artery segmentation, will be an important direction for future work.
V Conclusion
In this paper, we introduced VesselSAM, an enhanced version of the Segmentation Anything Model (SAM), designed specifically for aortic vessel segmentation. By incorporating the Atrous Attention Module and Low-Rank Adaptation (LoRA), VesselSAM addresses key limitations of the original SAM, enhancing its ability to capture complex hierarchical features typical of medical images. The Atrous Attention Module enables multi-scale feature extraction, effectively capturing both fine-grained details and broader anatomical context, while LoRA optimizes fine-tuning efficiency by reducing the number of trainable parameters without sacrificing performance.
Extensive experiments on the Aortic Vessel Tree (AVT) and Type-B Aortic Dissection (TBAD) datasets demonstrate that VesselSAM outperforms ViT-based and other SAM-based methods, achieving superior Dice Similarity Coefficient (DSC) and HD scores. The model achieves these results with significantly fewer trainable parameters, reinforcing its role as a parameter-efficient fine-tuning (PEFT) model for medical applications. These findings underline VesselSAM’s superior performance while maintaining computational efficiency, making it particularly valuable for real-world clinical tasks.
VesselSAM offers a robust solution for medical image segmentation, combining high accuracy with computational efficiency. Its ability to generalize across diverse datasets and perform well with minimal additional computational resources makes it highly suitable for clinical applications. Future work will explore further optimizations, including the integration of text-based prompts and visual language models, and expand its use to other medical imaging tasks, ensuring broader applicability in the healthcare domain.